OpenStudy (anonymous):

Please help! I give medals and become a fan! How do I multiply the two x-coordinates of the vectors and the two y-coordinates of v1=(22,15) and v2=(-2,4)?

OpenStudy (tkhunny):

Are you trying to calculate a dot product or inner product of some sort?

OpenStudy (anonymous):

v1 dot v2 = x1*x2 + y1*y2

OpenStudy (anonymous):

so 16?

OpenStudy (anonymous):

if you're after the dot product... yes

OpenStudy (jdoe0001):

\(\bf <a, b> \cdot <c, d> \implies a\times c + b\times d\)

OpenStudy (anonymous):

@tkhunny honestly I'm not really sure, it says write each product below and use the dot product to determine which of the following vector pairs are othogonal.

OpenStudy (jdoe0001):

hmm, didn't we cover this before?

OpenStudy (ybarrap):

you got it, but showing your work would make sure you weren't just lucky :)

OpenStudy (anonymous):

if the dot product is 0 then the vectors are orthogonal.

OpenStudy (ybarrap):

Do you know the definition of orthogonal?

OpenStudy (anonymous):

nope, what is it/

OpenStudy (jdoe0001):

lol

OpenStudy (ybarrap):

for example (1,0) and (0,1) are orthogonal? But why?

OpenStudy (ybarrap):

@pgpilot326 just presented the definition, but intuitively, do you know what it is?

OpenStudy (anonymous):

uhm when it equals zero?

OpenStudy (ybarrap):

For example, the y-axis itself it orthogonal to the x-axis, they are perpendicular to each other

OpenStudy (ybarrap):

Also, the angle between the vectors is 90 degrees. This comes from another way to write the dot product: \( |v1| |v2| cos \theta = v1*v2 \). If \(\theta\) is 90 degrees, then the dot product is zero.

OpenStudy (anonymous):

I need one more answer though

OpenStudy (anonymous):

nevermind I got it! Thanks everyone for your help!

OpenStudy (ybarrap):

np